# Template-Based Visual Program Distillation

### v2

@@ Appendix B Training and Model Details
Program generation settings for GPT can be found in Table 10. Template-based and direct training hyper-parameters can be found in Table 11. For CodeT5, the most important hyper-parameters were the learning rate and LoRA dropout rate. Training stopped when the training loss stopped decreasing.
All experiments were run on a single 40gb A40 or 40gb A100. Time measurements were measured on an A40.
@@
Table 11: Training and evaluation settings for CodeT5
Appendix C Changes to ViperGPT API
@@ Methods
Appendix E Variable Renamer
@@
Appendix F Variable Replacement
@@
Color red, blue, green, yellow, purple, black, white, orange, pink, brown, gray, indigo, cyan magenta, tan, silver
Activities running, walking, snowboarding, flying, splashing, tossing, riding, standing, hugging hanging, breaking, pulling, decorating, facing, preparing pouring, pointing, laughing
Relation picking up, in front of behind, above, below next to, near, far away close , following, on top, beside, walking on, attached, left, right, diagonal

### v4

@@ Appendix C Training and Model Details
Program generation settings for GPT can be found in Table 11. Training hyper-parameters can be found in Table 12. For distilled models, the most important hyper-parameters were the learning rate and LoRA dropout rate. Training stopped when the training loss stopped decreasing.
All experiments were run on a single 40gb A40 or 40gb A100. Time measurements were taken on an A40.
@@
Table 12: Training and evaluation settings for student models. We use the same learning rate for all models.
Appendix D Changes to ViperGPT API
@@ Methods
Appendix F Variable Renamer
@@
Appendix G Variable Replacement
@@
Color
red, blue, green, yellow,
purple, black, white,
orange, pink, brown,
gray, indigo, cyan
magenta, tan, silver
Activities
running, walking, snowboarding,
flying, splashing, tossing,
riding, standing, hugging
hanging, breaking, pulling,
decorating, facing, preparing
pouring, pointing, laughing
Relation
picking up, in front of
behind, above, below
next to, near, far away
close , following, on top,
beside, walking on, attached,
left, right, diagonal